40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16969     Entry time: Fri Jul 1 12:49:52 2022
Author: Koji 
Type: Update 
Category: IOO 
Subject: MC2 seemed misaligned / fixed 

I found the IMC was largely misaligned and was not locking. The WFS feedback signals were saturated and MC2 was still largely misaligned in yaw after resetting the saturation.
It seemed that the MC WFS started to put the large offset at 6:30AM~7:00AM (local).
MC2 was aligned and the lock was recovered then the MC WFS seems working for ~10min now.

Attachment 1: C1-MULTI_FBDB3F_TIMESERIES-1340668818-86400.png  62 kB  Uploaded Fri Jul 1 13:59:08 2022  | Hide | Hide all
C1-MULTI_FBDB3F_TIMESERIES-1340668818-86400.png
Attachment 2: C1-LOCKED_MC_5E4267_TIMESERIES-1340668818-86400.png  91 kB  Uploaded Fri Jul 1 13:59:13 2022  | Hide | Hide all
C1-LOCKED_MC_5E4267_TIMESERIES-1340668818-86400.png
ELOG V3.1.3-